Archive for October 2016

Apiary Record: 23rd October

A cloudy, windy and cold autumn day at only 10 C.   Only 5 members helped out putting mouse guards and woodpecker cages on all the hives.

Nuc 16 C  was combined, using the paper method, with Nuc 16 B, having first found and removed the Queen.     This was done on Thursday, a warmer day  ( 14-16 C ) and it was hoped to go through Sunday and rationalize the frames down to only one National Deep brood box.

We did lift the top brood box but the paper had only just been nibbled through and with the cold wind, it was thought prudent to delay the removal till a warmer day.

Next Apiary Meeting: This Sunday 23rd Oct. 14:00 Poss the last this year.

On Thursday, Graham and I combined the two Nuc’s , 16 C, with 16 B,  as the former was a small colony and was not likely to go through the winter.  It was also put onto a proper stand, next to the Nuc frame.     This Sunday we need to go through and rationalize the stores/brood, to leave a single deep over winter, the bees having eaten through the paper and combined by then.

Also on the “to do list” are the mouse guards and woodpecker cages.

Hope to see you all there for the last time this year.

Graham, Liz, Mel.  Apiary custodians..   and  Richard now a retired custodian.

Apiary Meeting: Today’s Meeting has been cancelled

Forecast for today is for possible heavy showers this afternoon, with gusty winds, so not worth the risk of disturbing the bees.    We will try again next week.

Next Apiary Meeting: Sunday 16th October 14:00, Weather dependent

Weather forecast is not looking good, so will update the website early Sunday morning to advise if on or off.      Most things can be left for a week, putting mouse guards and woodpecker cages on.     The most important job is to check through the Nuc 16C , as its a bit on the small side to go through the winter.  If it’s ok, and no disease, to combine it with the other Nuc,  16B.

Hope to see you there.

Richard, Liz, Mel and Graham Rowden :  Apiary Custodians

Apiary Record: 25th September 2016

Putting the colonies to bed for the winter, checking that they have enough stores and that entrances are down to a minimum.   Still a bit too early for mouse guards and woodpecker protection.    Removal of Varroa treatment and varroa boards also undertaken.

Colony #6   was on a National deep, but we combined Colony #1, which was on a 14×12 several weeks ago.   We had put the 14×12 under the deep after the two colonies were united, to encourage  them to take up the stores in it,  but they thought  otherwise and pulled out the entrance block, gaining access to the main entrance and so filled it with stores rather than taking it up into the National Deep. We went through the National Deep and used any stores left to help one of the Nuc’s and then removed it, leaving just the 14×12 and a full super on top.

We also decided to under-super the colonies,   putting full supers under the brood box’s, to help them get at the stores in the cold months to come.   Removing all the Queen excluders  to clean them ready for next season.

Weather was Sunny to start with  17 Degrees C ,  but became overcast with a shower of rain towards the end.  No wind to speak of.

Hive #2   Inspected by Richard  W.     Apiguard and Eke removed, top super 3/4 full and lots of stores in 14×12 Deep. Qx also removed.

Hive #3  Inspected by Richard W.     Apiguard and Eke removed, stores a bit light but incoming Balsam nectar and ivy could fill.  check next time..  may need feeding if weather bad.

Hive #4  Inspected by Richard W.  Apiguard and Eke removed.  Full super put under brood box and another full super put above a clearer board to give (bee free) to a nuc ( 16 B).

Hive #5  Inspected by Peter Reader,  Apiguard and Eke removed,  Full super put under brood box and Qx removed.

Hive #6  Inspected by Graham Rowden,  Apiguard and Eke removed, bees did not lift stores up into National Deep, so left in 14×12, ND removed, frames of stores given to 16 C.

Hive #7  Inspected by Richard W.   Apiguard and Eke removed. Qx removed and full super left on top as this has two ND brood boxes.

Hive #7A Inspected by Peter Reader. Apiguard removed with Eke. All Shallow boxes.  Qx removed,  top super full.

Hive #8  Inspected by Graham Rowden,   Apiguard and Eke removed.  Under-supered.

Hive #15F Inspected by Graham R,  Apiguard and Eke removed.

Nuc #16B  Inspected by Graham R.  Brood chamber – National Deep. Queen still laying nice brood pattern, BAS,  Full super from hive #4 added under deep.  colony to be used again for observation hive at Blackmoor.

Nuc #16C  Inspected by Graham R.   Brood chamber -National Deep, Light on stores, so given Deep frames from Hive#6,   Rapid feeder put on with 1 ltr sugar 1:1

No Apiary meeting this and next Sunday. Next meeting 16th October.

9th October  we are at Blackmoor  Apple Day, with a full display being organised by Graham Rowden.   Observation Hive, Honey Sales, Apple Cutting/tasting and lots of interesting displays. If you can help give Graham a call.

Next full meeting is on Sunday 16th October when we will be Hefting hives, checking stores ect.